# Finely Stratified Rerandomization Designs

### v2

@@ 4 Nonlinear Rerandomization
In this section, we study several novel “nonlinear” rerandomization criteria, proving that in many cases such criteria are asymptotically equivalent to linear rerandomization (Definition 2.1), with an implicit choice of rerandomization covariates $h$ and acceptance region $A$. This shows that our asymptotics and inference methods apply to a broad class of asymptotically linear rerandomization schemes.
@@ 4.1 GMM Rerandomization
First, we generalize the imbalance metric $\mathcal{I}_{n}$ introduced in Definition 2.1, allowing rejection of a treatment allocation $D_{1:n}$ based on potentially nonlinear features of the in-sample distribution of treatments and covariates $(D_{i},X_{i})_{i=1}^{n}$. We can define a large class of nonlinear imbalance metrics by letting $m(X_{i},\beta)$ be a score function and considering within-arm GMM estimators $\widehat{\beta}_{1}$ and $\widehat{\beta}_{0}$ defined by
@@
We propose to rerandomize until the within-arm parameter estimates are approximately equal, $\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\approx 0$.
@@
Intuitively, the generalized imbalance metric $\mathcal{I}_{n}^{m}$ allows us to randomize until possibly nonlinear features of the covariates are balanced between the treatment and control groups. Observe that if $m(X_{i},\beta)=X_{i}-\beta$, then $\widehat{\beta}_{d}=\bar{X}_{d}$ for $d=0,1$ and $\mathcal{I}_{n}^{m}=\mathcal{I}_{n}$, so linear rerandomization is a special case.
@@
Let $f(X,\beta)$ be a parametric density model for covariates $X$, which may be misspecified. After drawing $D_{1:n}$ by stratified randomization, consider forming (quasi) maximum likelihood estimators $\widehat{\beta}_{1}\in\operatorname*{argmax}_{\beta}E_{n}[D_{i}\log f(X_{i},\beta)]$ and $\widehat{\beta}_{0}\in\operatorname*{argmax}_{\beta}E_{n}[(1-D_{i})\log f(X_{i},\beta)]$ for the density of covariates assigned to each treatment arm, rerandomizing until the estimated parameters $\sqrt{n}|\widehat{\beta}_{1}-\widehat{\beta}_{0}|_{2}\leq\epsilon$. Under regularity conditions, [^17] For example, if $\beta\to\log f(X,\beta)$ is a.s. strictly concave, the key identification condition in Assumption 3.2 will be satisfied. $\widehat{\beta}_{d}$ are GMM estimators as in Equation 4.1 with score function $m(X_{i},\beta)=\nabla_{\beta}\log f(X_{i},\beta)$, so this procedure is a GMM rerandomization with acceptance region $A=\{x:|x|_{2}\leq\epsilon\}$.
Let $\beta^{*}$ be the unique solution to $E[m(X,\beta^{*})]=0$ and define $G_{m}=E[(\partial/\partial\beta^{\prime})m(X_{i},\beta^{*})]$. Our next result shows that GMM rerandomization with acceptance criterion $\mathcal{I}_{n}^{m}\in A$ is equivalent to linear rerandomization (Definition 2.1) with an implicit choice of rerandomization covariates $h_{i}=m(X_{i},\beta^{*})$ and linearly transformed acceptance region.
@@
$V_{a}=\min_{\gamma\in\mathbb{R}^{d_{m}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\gamma^{\prime}m(X_{i},\beta^{*})|\psi)].$ (4.2)
The residual $R\sim\gamma_{0}^{\prime}Z_{m}\,|\,Z_{m}\in G_{m}A$ for $Z_{m}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(m(X_{i},\beta^{*})|\psi)])$, where $\gamma_{0}$ is optimal in Equation 4.2.
Theorem 4.3 shows that by rerandomizing until $\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\in A$, we implicitly balance the influence function $-G_{m}^{-1}m(X_{i},\beta^{*})$ for the difference of GMM estimators in Equation 4.1. This suggests an equivalent, but computationally much simpler design with only one round of nonlinear estimation. In particular, let $\widehat{\beta}$ solve $E_{n}[m(X_{i},\widehat{\beta})]=0$ be the pooled GMM estimator and set rerandomization covariates $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$, rerandomizing until $\sqrt{n}E_{n}[H_{i}\widehat{h}_{i}]\in G_{m}A$. The next result shows that this design, which generalizes Definition 2.1 to allow for estimated covariates, is asymptotically equivalent to the GMM rerandomization in Definition 4.1.
Corollary 4.4.
Suppose Assumption 3.1, 3.2 hold and let $m(X,\beta)$ be as in Definition 4.1. Let $D_{1:n}$ be rerandomized as in Definition 2.1 with $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$ and acceptance region $G_{m}A$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R$, with both variables identical to those in Theorem 4.3.
Corollary 4.4 is a useful tool for showing the equivalence of computationally intensive designs based on nonlinear estimation with simpler linear schemes. For instance, Example 4.5 shows that density rerandomization over an exponential family with sufficient statistic $r(X_{i})$ is asymptotically equivalent to linear rerandomization with $h_{i}=r(X_{i})$.
@@
The sufficient statistic is $r(x)=((x_{j})_{j},(x_{j}x_{l})_{j<l})$, and the parameters $\beta_{jl}$ model correlation between the binary variables $x_{j}$ and $x_{l}$. By the discussion above, a design that rerandomizes based on the difference of quasi-MLE density estimates in this family [^20] This is well-motivated when $\psi$ is expected to be more important than $(x_{j})_{j}$. We don’t want to stratify on both, since this could radically decrease match quality on $\psi$. is asymptotically equivalent to the much simpler linear rerandomization in Definition 2.1 with covariates $h_{i}=((x_{j})_{j},(x_{j}x_{l})_{j<l})$.
@@ 4.2 Propensity Score Rerandomization
To motivate a propensity score based rerandomization procedure, note that under stratified randomization we have $E[D_{i}|X_{i}]=p$ for all units. In finite samples, however, the realized propensity $\widehat{p}(B)=E_{n}[D_{i}|X_{i}\in B]$ may significantly diverge from $p$ in certain regions $B\subseteq\mathbb{R}^{d_{X}}$ of the covariate space. This implies that covariates are predictive of treatment assignments post-randomization, a form of “in-sample confounding,” which of vanishes as $n\to\infty$ but affects precision. To prevent this, we could, for instance, reject allocations where $|\widehat{p}(B)-p|>\epsilon$ for some collection of sets $B$. To make this idea tractable without fully discretizing, consider a parametric propensity model $p(X,\beta)=L(X^{\prime}\beta)$ and define the MLE estimator
@@
We can measure the average gap between the estimated and true propensity score using
@@
Do the following: (1) form groups as in Definition 2.1. (2) Draw $D_{1:n}$ and estimate the propensity model in Equation 4.3. (3) If imbalance $\mathcal{J}_{n}\leq\epsilon$, accept. Otherwise, repeat from (2).
@@
Figure 1: Propensity rerandomization (Definition 4.6) with $p=1/2$ for $Z\sim\operatorname{Unif}[0,1]$ and $X=B(Z)$ a B-spline basis. LHS: $D_{1:n}$ and estimated propensity with $\widehat{p}(Z)\ll 1/2$, for $Z\in[0.4,0.9]$. RHS: Accepted allocation $D_{1:n}$ with $\mathcal{J}_{n}\leq\epsilon$
Our next result shows that propensity rerandomization as in Definition 4.6 is equivalent to a simpler linear rerandomization design, with an implicit choice of ellipsoidal acceptance region. We require some extra regularity conditions on the link function $L$, which for brevity we state in Appendix 8.5.
@@
Suppose $D_{1:n}$ is as in Definition 4.6. Require Assumptions 3.2, 8.12. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R$.
@@
The residual $R\sim\gamma_{0}^{\prime}Z_{h}\,|\,Z_{h}^{\prime}\operatorname{Var}(h)^{-1}Z_{h}\leq\epsilon$ for $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$ and $\gamma_{0}$ optimal in the equation above.
Theorem 4.7 shows that for any sufficiently regular link function, propensity rerandomization is asymptotically equivalent to the quadratic rerandomization design in Example 2.2, with acceptance criterion $n(\bar{h}_{1}-\bar{h}_{0})^{\prime}\operatorname{Var}_{n}(h_{i})^{-1}(\bar{h}_{1}-\bar{h}_{0})\leq\epsilon$. Equivalently, propensity rerandomization behaves like linear rerandomization with $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ and ellipsoidal acceptance region $A=\operatorname{Var}(h)^{1/2}B(0,\epsilon)$. [^21] A related result was found by Ding and Zhao, (2024), who study rerandomizing until the p-value of a logistic regression coefficient is above a threshold.
Implicit Acceptance Regions. Both nonlinear designs in this section turned out to be equivalent to the standard rerandomization scheme in Definition 2.1, with a specific, implicit choice of rerandomization moments and acceptance region determined by the choice of score $m$ and marginal covariate distribution. However, this implicit choice is not likely to be optimal, since the residual term in the asymptotic error distribution $R_{A}\sim\gamma_{0}^{\prime}Z_{h}|Z_{h}\in A$ depends on both the covariates $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$, and the partially linear coefficient $\gamma_{0}$. This coefficient is determined by the joint distribution of the assignment influence function $\Pi a(W,\theta_{0})$ and covariates $(\psi,h)$. In the next section, we show how to use prior information about this joint distribution to optimize the acceptance region and bound the variance of $R_{A}$.
@@ 5 Optimizing Acceptance Regions
In this section, we study efficient choice of the acceptance region $A\subseteq\mathbb{R}^{d_{h}}$. We propose a novel minimax rerandomization criterion that minimizes asymptotic computational cost subject to a strict lower bound on statistical efficiency.
For intuition, we first restrict to the case of estimating $\theta_{n}=\operatorname{SATE}$, generalizing in what follows. Example 3.9 showed that $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow\mathcal{N}(0,V(\gamma_{0}))+\gamma_{0}^{\prime}Z_{hA}$, independent RV’s with $Z_{hA}=Z_{h}|Z_{h}\in A$ and $V(\gamma_{0})$ that does not depend on $A$. The coefficient $\gamma_{0}$ is identified by the partially linear regression [^22] This expansion is without loss of generality. We do not impose well-specification $E[e|\psi,h]=0$.
$\bar{Y}=\gamma_{0}^{\prime}h+t_{0}(\psi)+e,\quad\;E[e|\psi]=0,\;E[eh]=0.$
The term $R_{A}=\gamma_{0}^{\prime}Z_{hA}$ arises from fluctuations of $\bar{Y}$ predictable by $h$ that are not balanced by rerandomization due to slackness in the acceptance region, $A\not=\{0\}$. More precisely, $\gamma_{0}^{\prime}Z_{hA}$ is the rerandomized limiting distribution of $\gamma_{0}^{\prime}\mathcal{I}_{n}=\gamma_{0}^{\prime}\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$, the projection of covariate imbalances in $h$ along the direction $\gamma_{0}$. This suggests an oracle acceptance criterion that rerandomizes until $|\gamma_{0}^{\prime}\mathcal{I}_{n}|\leq\epsilon$, with acceptance region $A=\{x:|\gamma_{0}^{\prime}x|\leq\epsilon\}$. Of course, this is infeasible since $\gamma_{0}$ is unknown when designing the experiment.

### v3

@@ 4 Nonlinear Rerandomization
In this section, we introduce several novel “nonlinear” rerandomization criteria, proving that in many cases such designs are first-order equivalent to linear rerandomization (Definition 2.1), with an implicit choice of covariates $h$ and acceptance region $A$. This shows that our asymptotics and inference methods apply to a broad class of asymptotically linear rerandomization schemes, expanding the scope of the results in Section 3 above.
@@ 4.1 GMM Rerandomization
First, we generalize the imbalance metric $\mathcal{I}_{n}$ in Definition 2.1, allowing rejection of $D_{1:n}$ based on potentially nonlinear features of the in-sample distribution of treatments and covariates $(D_{i},X_{i})_{i=1}^{n}$. Let $m(X_{i},\beta)$ be a GMM score function, separate from the score $g$ defining the estimands above. We can define a large class of interesting designs by stratifying and rerandomizing until $\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\approx 0$ for within-arm GMM estimators
@@
Observe that if $m(X_{i},\beta)=X_{i}-\beta$, then $\widehat{\beta}_{d}=\bar{X}_{d}$ for $d=0,1$ and $\mathcal{I}_{n}^{m}=\mathcal{I}_{n}$, so linear rerandomization is a special case. However, Definition 4.1 also allows for novel designs, such as rerandomizing until the estimated densities of covariates $X_{i}|D_{i}=1$ among treated and $X_{i}|D_{i}=0$ among control are similar. To the best of our knowledge, we provide the first formal results for such a design.
@@
Let $f(X,\beta)$ be a possibly misspecified parametric density model for covariates $X$. After drawing $D_{1:n}$ by stratified randomization, consider forming (quasi) maximum likelihood estimators $\widehat{\beta}_{1}\in\operatorname*{argmax}_{\beta}E_{n}[D_{i}\log f(X_{i},\beta)]$ and $\widehat{\beta}_{0}\in\operatorname*{argmax}_{\beta}E_{n}[(1-D_{i})\log f(X_{i},\beta)]$, rerandomizing until the estimated parameters $\sqrt{n}|\widehat{\beta}_{1}-\widehat{\beta}_{0}|_{2}\leq\epsilon$. Under regularity conditions, [^20] For example, if $\beta\to\log f(X,\beta)$ is a.s. strictly concave, the key identification condition in Assumption 3.2 will be satisfied. $\widehat{\beta}_{d}$ are GMM estimators as in Equation 4.1 with score function $m(X_{i},\beta)=\nabla_{\beta}\log f(X_{i},\beta)$, so this procedure is a GMM rerandomization with acceptance region $A=\{x:|x|_{2}\leq\epsilon\}$.
Let $\beta^{*}$ be the unique solution to $E[m(X,\beta^{*})]=0$ and define $G_{m}=E[(\partial/\partial\beta^{\prime})m(X_{i},\beta^{*})]$. Our next result shows that GMM rerandomization with acceptance criterion $\mathcal{I}_{n}^{m}\in A$ is equivalent to linear rerandomization (Definition 2.1) with an implicit choice of rerandomization covariates $h_{i}=m_{i}^{*}\equiv m(X_{i},\beta^{*})$ and linearly transformed acceptance region.
@@
$V_{a}=\min_{\gamma\in\mathbb{R}^{d_{m}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\gamma^{\prime}m_{i}^{*}|\psi)].$ (4.2)
The residual $R\sim[\gamma_{0}^{\prime}Z_{m}\,|\,Z_{m}\in G_{m}A]$ for $Z_{m}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(m_{i}^{*}|\psi)])$, where $\gamma_{0}$ is optimal in Equation 4.2.
Theorem 4.3 shows that by rerandomizing until $\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\in A$, we implicitly balance the influence function $-G_{m}^{-1}m(X_{i},\beta^{*})$ for the difference of GMM estimators above. In particular, this shows that all GMM rerandomization designs are first-order equivalent to linear rerandomization (Definition 2.1) for some choice of $h_{i}$ and acceptance region $A$.
[…]
Corollary 4.4 (Feasible Equivalence).
Suppose Assumption 3.1, 3.2 and let $m(X,\beta)$ as in Definition 4.1. Let $D_{1:n}$ be rerandomized as in Definition 2.1 with $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$ and acceptance region $\widehat{G}_{m}A$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R$, with both variables identical to those in Theorem 4.3.
One consequence of Corollary 4.4 is that density based rerandomization for likelihood in an exponential family with sufficient statistic $r(X_{i})$ is asymptotically equivalent to linear rerandomization setting $h_{i}=r(X_{i})$.
@@
This is an exponential family with sufficient statistic $r(x)=((x_{j})_{j},(x_{j}x_{l})_{j<l})$. The parameters $\beta_{jl}$ model correlation between the binary variables $x_{j}$ and $x_{l}$. For $x\in\{\pm 1\}^{k}$ with $k$ large, this is a tractable alternative to nonparametricallly modeling the full joint distribution, or e.g. stratifying on all $2^{k}$ cells. Corollary 4.4 shows that rerandomizing based on the difference of quasi-MLE density estimates in this family [^23] This is well-motivated when $\psi$ is expected to be more important than $(x_{j})_{j}$. We don’t want to stratify on both, since this could radically decrease match quality on $\psi$. is asymptotically equivalent to a simpler linear rerandomization design with $h_{i}=((x_{j})_{j},(x_{j}x_{l})_{j<l})$.
@@ 4.2 Propensity Score Rerandomization
To motivate a propensity score based rerandomization procedure, note that despite $E[D_{i}|X_{i}]=p$ for all units, in finite samples the realized propensity $\widehat{p}(B)=E_{n}[D_{i}|X_{i}\in B]$ may significantly diverge from $p$ in certain regions $B\subseteq\mathbb{R}^{d_{X}}$ of the covariate space. This implies that covariates $X_{i}$ are predictive of treatment assignments $D_{i}$ ex-post, a form of “in-sample confounding,” which vanishes as $n\to\infty$ but affects precision. To prevent this, we could reject allocations where $|\widehat{p}(B)-p|>\epsilon$ for some collection of sets $B$. To make this idea tractable without fully discretizing, consider a parametric propensity model $p(X,\beta)=L(X^{\prime}\beta)$ for smooth link function $L$ (e.g. Logit) and define the MLE estimator
@@
The average gap between the realized and ex-ante propensity score can be measured by
@@
Do the following: (1) form groups as in Definition 2.1. (2) Draw $D_{1:n}$ by stratified randomization and estimate the propensity model in Equation 4.3. (3) If imbalance $\mathcal{J}_{n}\leq\epsilon^{2}$, accept. Otherwise, repeat from (2).
@@
Figure 1: Propensity rerandomization (Definition 4.6) with $p=1/2$ for $Z\sim\operatorname{Unif}[0,1]$ and $X=B(Z)$ a B-spline basis. LHS: $D_{1:n}$ and estimated propensity with $\widehat{p}(Z)\ll 1/2$, for $Z\in[0.4,0.9]$, showing imbalance. RHS: Accepted allocation $D_{1:n}$ with $\mathcal{J}_{n}\leq\epsilon$
This design is illustrated in Figure 1. Note that the covariate distribution is approximately balanced between $D=1$ and $D=0$ after acceptance. Our next result shows that propensity rerandomization as in Definition 4.6 is equivalent to a simpler linear rerandomization design, with an implicit choice of ellipsoidal acceptance region. We require some extra regularity conditions on the link function $L$, which for brevity we state in Appendix 12.5.
@@
Suppose $D_{1:n}$ is as in Definition 4.6. Require Assumptions 3.2, 12.12. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R$.
@@
The residual $R\sim\gamma_{0}^{\prime}Z_{h}\,|\,Z_{h}^{\prime}\operatorname{Var}(h)^{-1}Z_{h}\leq\epsilon v_{D}^{-2}$ for $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$ and $\gamma_{0}$ optimal in the equation above.
Theorem 4.7 shows that for any sufficiently regular link function, [^24] Theorem 4.7 uses MLE estimation of $\widehat{\beta}$, though we conjecture the result would be identical for inverse probability tilting (Graham, (2012)) or tailored loss function (Zhao, (2019)) estimation. propensity rerandomization is asymptotically equivalent to Mahalanobis rerandomization in Example 2.4, with acceptance criterion $n(\bar{h}_{1}-\bar{h}_{0})^{\prime}\operatorname{Var}_{n}(h_{i})^{-1}(\bar{h}_{1}-\bar{h}_{0})\leq\epsilon v_{D}^{-2}$. Equivalently, propensity rerandomization behaves like linear rerandomization with $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ and ellipsoidal acceptance region $A=\operatorname{Var}(h)^{1/2}B(0,\epsilon v_{D}^{-2})$. [^25] A related result was found by Ding and Zhao, (2024), who study rerandomizing until the p-value of a logistic regression coefficient is above a threshold.
This section introduced a large family of novel rerandomization methods based on nonlinear estimators. Theorems 4.3 and 4.7 broaden the scope of our asymptotic theory and inference results, showing they also apply to these designs. These equivalence results raise the bar for future methodology improvements, showing that to obtain rerandomization designs with different first-order properties, we may need to consider more stringent imbalance measures, such as nonparametric two-sample test statistics. We leave this extension to future work.
Motivated by the “implicit” acceptance regions chosen by the designs in this section, next we formally study optimal choice of the acceptance region $A$.
@@ 5 Optimizing Acceptance Regions
In this section, we study efficient choice of the acceptance region $A\subseteq\mathbb{R}^{d_{h}}$. We propose a novel minimax rerandomization scheme and show that it minimizes the computational cost of rerandomization subject to a strict lower bound on statistical efficiency. This can be viewed as a form of dimension reduction, increasing rerandomization acceptance probability by downweighting less important directions in the covariate space $h$.
For simplicity, we first restrict to the case of estimating $\theta_{n}=\operatorname{SATE}$. Example 3.8 showed that $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow\mathcal{N}(0,V(\gamma_{0}))+\gamma_{0}^{\prime}Z_{hA}$, independent RV’s with $Z_{hA}=Z_{h}|Z_{h}\in A$ and variance $V(\gamma_{0})$ that does not depend on $A$. The term $Z_{hA}$ arises from residual imbalances in $h$ due to slackness in the acceptance region, $A\not=\{0\}$. The coefficient $\gamma_{0}$ comes from the partially linear regression [^26] This expansion is without loss of generality. We do not impose well-specification $E[e|\psi,h]=0$.
$\bar{Y}=\gamma_{0}^{\prime}h+t_{0}(\psi)+e,\quad\;E[e|\psi]=0,\;E[eh]=0.$ (5.1)
All together, the residual imbalance term $\gamma_{0}^{\prime}Z_{hA}$ is the limiting distribution under rerandomization of $\gamma_{0}^{\prime}\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$, the projection of covariate imbalances in $h$ along the direction $\gamma_{0}$. This suggests an oracle acceptance criterion that rerandomizes until the imbalance $|\gamma_{0}^{\prime}\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})|\leq\epsilon$, with acceptance region $A=\{x:|\gamma_{0}^{\prime}x|\leq\epsilon\}$, reducing the problem to one dimension from arbitrary $\dim(h)$. Of course, this oracle design is infeasible since $\gamma_{0}$ is generally unknown when designing the experiment.